# Durham, UK

### v1

@@ 5 Summary
There are significant differences between the two scenarios. If light ALPs are dark matter, several effects can be observed that are not present for a light ALP field that does not contribute significantly to the dark matter relic density. The variation of fundamental constants induced by oscillations of the quadratic ALP field is the most sensitive probe for ALPs lighter than $m_{a}\approx 10^{-20}$ eV. These experiments probe the ALP gluon coupling via its correction to nucleon masses, photons, and electrons. Quadratic ALP interactions probe the gradient of the ALP field which leads to very strong constraints from tests of the equivalence principle for masses $m_{a}\approx 10^{-20}-10^{-5}$ eV. We stress however that the assumptions going into the calculation of both variations of fundamental constants and tests of the equivalence principle rely on a small-coupling approximation which breaks down for an interaction strength of $c_{GG}/f\gtrsim 10^{-15}$ GeV-1, as indicated by the cyan dashed line in Figure 8. For larger couplings, any quadratically interacting dark matter field has nonlinear solutions which in the case of the ALP leads to an unphysical parameter space shortpaper . Haloscopes and helioscopes mostly probe the ALP photon coupling via the Primakoff effect in an external magnetic field. Even though haloscopes probe the linear ALP-photon interaction, they measure the local ALP field which is affected by the non-linear behaviour induced by quadratic interactions. In the absence of this effect, they are the most sensitive experiments in the mass range $m_{a}\approx 10^{-5}-10^{-4}$ eV. Helioscopes are insensitive to the dark matter halo and the corresponding parameter space can be considered ruled out independent of the local value of the ALP field.
In the case where the ALP field does not contribute to the dark matter, the experimental sensitivity is not affected by the nonlinear field values of dark matter ALPs close to massive bodies. The strongest constraints set by lab measurements for the whole mass range in this case are set by the CAST helioscope. We also show the parameter space excluded by searches for fifth forces which include the contributions from shift-invariance breaking quadratic ALP-nucleon interaction. Both probe the couplings induced by the RG running.
Figure 8: Current experimental limits (sensitivity) on ALP dark matter from - (i) quantum sensors (ii) haloscopes and (iii) EP-violation tests. Exclusions are shown in solid lines for atomic clocks in red, clock-cavity tests in blue, optical/laser interferometers and mechanical resonators in green, haloscopes in orange and EP-violation tests in black. Also shown are projections in dashed lines from nuclear clock (Th-229) in red, atom interferometers (broadband and resonant modes) in purple, mechanical resonators in green, haloscopes in orange and nuclear magnetic resonances in magenta. Above the cyan dashed line, the small-coupling approximation breaks down, leading to an unphysical parameter space.
Figure 9: Current experimental limits (sensitivity) on the ALP-gluon coupling at the UV scale where we compare only the most stringent bounds for each mass from two categories - (i) helioscopes in blue and (ii) fifth forces in red, orange and brown solid lines and shades. The future projections are shown with dashed lines. In helioscopes, we show the current exclusions from CAST and future projections from IAXO. The current limits for fifth-forces are taken from Rb -21Ne comagnetometer, torsion balances and Casimir-less forces. All these experiments do not require the ALP to be dark matter.
@@ 6 Conclusions
We present a comprehensive analysis of the sensitivity of quantum sensors and high-precision measurements to effects from light axions or axionlike particles. Below a mass of a few electronvolt ALPs can be dark matter candidates, with the relic density produced via misalignment and the ALP field behaves like a classical (pseudo)scalar background field. We compute the most promising experimental strategy for light ALP searches and stress the complementarity of experiments that can distinguish whether ALPs make up a substantial component of dark matter.
@@
To this end, we go beyond the existing work on quantum sensor searches for ALPs in several important ways. First, we take effects from renormalisation group running and matching into account by expressing the low energy couplings of the ALP with SM fields in terms of the coefficients of the UV theory, which allows to compare the sensitivity of experiments at different scales. An important result is the ALP coupling to electrons that is generated via RG effects present even if the ALP only couples to gluons in the UV. We further use the ALP interactions in the chiral Lagrangian to compute interactions quadratic in the ALP field. We also express the ‘dilatonic charges’ in terms of the ALP couplings in the chiral Lagrangian. These interactions are nominally subleading with respect to linear ALP interactions, but induce effects that can be more relevant due to the experimental sensitivity, for example, time-dependent fundamental constants and spin-independent fifth forces. We find a clear hierarchy in quadratic ALP couplings due to the derivative nature of ALP couplings, which always leads to a suppression of quadratic ALP-electron couplings. The solution of the equations of motion for quadratically coupled fields in the presence of a massive body such as a planet has a position-dependent mass term which implies an unphysical parameter space for ALPs for $c_{GG}/f\gtrsim 10^{-15}$ GeV-1. We emphasize that experiments that are sensitive to ALPs independent of whether they are dark matter such as helioscopes and searches for fifth forces can constrain this parameter space, whereas the calculations for others that are sensitive to the local ALP field value are not consistent.
@@
Even though our formalism allows to consider any combination of ALP couplings in the UV, we use a scenario with a single ALP coupling to gluons in the UV in order to illustrate our results. We compare constraints from the fifth forces induced by linear and quadratic ALP interactions and emphasize how quadratic interactions can lead to better sensitivity since linear ALP couplings only induce spin-dependent forces. The results are compared with constraints from Big Bang nucleosynthesis which is sensitive to variations in the neutron-proton mass splitting. The sensitivity of cavity searches with haloscopes and helioscopes is also significantly modified by taking into account quadratic ALP interactions. Besides a resonance induced by the conversion of a single pseudoscalar in the external magnetic field, cavities are also sensitive to the scalar interaction of two ALP fields which put constraints in a different mass range. Quantum clocks are one of the most sensitive probes of very light ALPs by looking for variations in fundamental constants. We compare the sensitivity of optical, microwave, future nuclear, and molecular clocks with cavity searches. For low masses, clocks are the most sensitive experiments, whereas tests of the equivalence principle and haloscopes are more sensitive for higher masses. For intermediate masses, mechanical resonators are competitive with haloscopes.
@@
There are several directions for future work. A full solution of the field equations close to massive bodies, taking into account the backreaction from the metric, is necessary to compute reliable theory predictions for the unphysical parameter space for almost all ALP scenarios. Different UV models, in particular models with photon and lepton couplings, would substantially change the hierarchy of experimental limits and projections. Finally, taking into account CP violation in the Standard Model gives rise to ALP-induced neutron dipole moments, spin precession, and long-range forces from a single ALP exchange, which could considerably affect the allowed parameter space.
@@ 7 Acknowledgements
MB thanks Joerg Jaeckel and Michael Spannowsky for useful discussions. MB and SC acknowledge support from the UKRI Future Leader Fellowship DarkMAP.

### v2

@@ 4.5 Atom Interferometers
Figure 8: Current experimental limits (sensitivity) on ALP dark matter from - (i) quantum sensors (ii) haloscopes and (iii) EP-violation tests. Exclusions are shown in solid lines for atomic clocks in red, clock-cavity tests in blue, optical/laser interferometers and mechanical resonators in green, haloscopes in orange and EP-violation tests in brown. Also shown are projections in dashed lines from nuclear clock (Th-229) in red, atom interferometers (broadband and resonant modes) in purple, mechanical resonators in green, haloscopes in orange and nuclear magnetic resonances in magenta. Above the black dashed line, the small-coupling approximation breaks down, leading to an unphysical parameter space.
Figure 9: Current experimental limits (sensitivity) on the ALP-gluon coupling at the UV scale where we compare only the most stringent bounds for each mass from two categories - (i) helioscopes in blue and (ii) fifth forces in red, orange and brown solid lines and shades. The future projections are shown with dashed lines. In helioscopes, we show the current exclusions from CAST and future projections from IAXO. The current limits for fifth-forces are taken from Rb -21Ne comagnetometer, torsion balances and Casimir-less forces. All these experiments do not require the ALP to be dark matter. The bound from non-perturbative ALP field values close to earth is shown by the green dashed contour, see  Hook:2017psm
@@ 5 Summary
There are significant differences between the two scenarios. If light ALPs are dark matter, several effects can be observed that are not present for a light ALP field that does not contribute significantly to the dark matter relic density. The variation of fundamental constants induced by oscillations of the quadratic ALP field is the most sensitive probe for ALPs lighter than $m_{a}\approx 10^{-20}$ eV. These experiments probe the ALP gluon coupling via its correction to nucleon masses, photons, and electrons. Quadratic ALP interactions probe the gradient of the ALP field which leads to very strong constraints from tests of the equivalence principle for masses $m_{a}\approx 10^{-20}-10^{-5}$ eV. We stress however that the assumptions going into the calculation of both variations of fundamental constants and tests of the equivalence principle rely on a small-coupling approximation which breaks down for an interaction strength of $c_{GG}/f\gtrsim 10^{-15}$ GeV-1 and $m_{a}\lesssim 7\times 10^{8}/f\ {\rm eV^{2}}$, as indicated by the black dashed line in Figure 8. For larger couplings, the ALP dark matter field has nonlinear solutions which leads to an non-perturbative parameter space shortpaper . Haloscopes and helioscopes mostly probe the ALP photon coupling via the Primakoff effect in an external magnetic field. Even though haloscopes probe the linear ALP-photon interaction, they measure the local ALP field which is affected by the non-linear behaviour induced by quadratic interactions. In the absence of this effect, they are the most sensitive experiments in the mass range $m_{a}\approx 10^{-5}-10^{-4}$ eV. Helioscopes are insensitive to the dark matter halo and the corresponding parameter space can be considered ruled out independent of the local value of the ALP field.
In the case where the ALP field does not contribute to the dark matter, non-perturbative ALP field values are relevant close to massive bodies for small ALP masses Hook:2017psm . The strongest constraints set by lab measurements for the whole mass range in this case are set by the CAST helioscope. We also show the parameter space excluded by searches for fifth forces which include the contributions from shift-invariance breaking quadratic ALP-nucleon interaction. Both probe the couplings induced by the RG running.
@@ 6 Conclusions
We present a comprehensive analysis of the sensitivity of quantum sensors and high-precision measurements to effects from light axions or axionlike particles. Below a mass of a few electronvolt ALPs can be dark matter candidates, with the relic density produced via misalignment and the ALP field behaving like a classical (pseudo)scalar background field. We compute the most promising experimental strategy for light ALP searches and stress the complementarity of experiments that can distinguish whether ALPs make up a substantial component of dark matter.
@@
To this end, we go beyond the existing work on quantum sensor searches for ALPs in several important ways. First, we take effects from renormalisation group running and matching into account by expressing the low energy couplings of the ALP with SM fields in terms of the coefficients of the UV theory, which allows to compare the sensitivity of experiments at different scales. An important effect is the ALP coupling to electrons that is generated via RG effects present even if the ALP only couples to gluons in the UV. We further use the ALP interactions in the chiral Lagrangian to compute interactions quadratic in the ALP field. We also express the ‘dilatonic charges’ in terms of the ALP couplings in the chiral Lagrangian. These interactions are nominally subleading with respect to linear ALP interactions, but induce effects that can be more relevant due to the experimental sensitivity, for example, time-dependent fundamental constants and spin-independent fifth forces. We find a clear hierarchy in quadratic ALP couplings due to the derivative nature of ALP couplings, which always leads to a suppression of quadratic ALP-electron couplings. The solution of the equations of motion for quadratically coupled fields in the presence of a massive body such as a planet has a position-dependent mass term which implies an non-perturbative field values for ALPs for $c_{GG}/f\gtrsim 10^{-15}$ GeV-1 and $m_{a}\lesssim 7\times 10^{8}/f\ {\rm eV^{2}}$, an effect equivalent to the one observed for ALPs that are not dark matter in Hook:2017psm .
@@
Even though our formalism allows to consider any combination of ALP couplings in the UV, we use a scenario with a single ALP coupling to gluons in the UV in order to illustrate our results. We compare constraints from the fifth forces induced by linear and quadratic ALP interactions and emphasize how quadratic interactions can lead to better sensitivity since linear ALP couplings only induce spin-dependent forces. The sensitivity of cavity searches with haloscopes and helioscopes is also significantly modified by taking into account quadratic ALP interactions. Besides a resonance induced by the conversion of a single pseudoscalar in the external magnetic field, cavities are also sensitive to the scalar interaction of two ALP fields which put constraints in a different mass range. Quantum clocks are one of the most sensitive probes of very light ALPs by probing variations in fundamental constants. We compare the sensitivity of optical, microwave, future nuclear, and molecular clocks with cavity searches. For low masses, clocks are the most sensitive experiments, whereas tests of the equivalence principle and haloscopes are more sensitive for higher masses. For intermediate masses, mechanical resonators are competitive with haloscopes.
@@
There are several directions for future work. A full solution of the field equations close to massive bodies, taking into account the dark matter boundary conditions and the full ALP potential, is necessary to compute reliable theory predictions for the non-perturbative parameter space. Different UV models, in particular models with photon and lepton couplings, would substantially change the hierarchy of experimental limits and projections. Finally, taking into account CP violation in the Standard Model gives rise to ALP-induced neutron dipole moments, spin precession, and long-range forces from a single ALP exchange, which could considerably affect the allowed parameter space.
@@ 7 Acknowledgements
MB thanks Joerg Jaeckel and Michael Spannowsky for useful discussions. We further thank Itai Bloch, Quentin Bonnefoy, Kai Bartnick, Sebastian Ellis, Gilad Perez, Inbar Savoray, Javi Serra, Konstantin Springmann, Stefan Stelzl and Andreas Weiler for useful comments and in particular for pointing out work on non-perturbative axion field configurations in the context of astrophysical observables. MB and SC acknowledge support from the UKRI Future Leader Fellowship DarkMAP.